# On integer distance sets

### v2

@@ 5.1. Lines
Proposition 1.4 was essentially proved in [24, Section 5]. However, since some of the terms used in [24] are nonstandard and [24, Theorem 4] is phrased in a manner that is not equivalent to Proposition 1.4, we will present a detailed proof of Proposition 1.4 here.
@@
|$\left|\{Db_{i}+D(q_{1}-m_{i})\colon i=1,\dots,K\}\right|\asymp K.$
@@
using the standard estimate $\tau(n)\ll n^{O(1/\log\log n)}$ (see, for example, [17, Theorem 315]), which gives the claim. ∎
@@ 5.2. Circles
Recall that Proposition 1.5 controls the size of an integer distance set in $[-N,N]^{2}$ that lives on an arbitrary circle. It turns out that the bulk of the argument lies in tackling the case where the circle is fully contained in $[-N,N]^{2}$; the general case easily follows. In particular, for any positive real number $r$, define $c(r)$ to be the maximum size of an integer distance set contained in the circle of radius $r$ centered at the origin. Then, this special case of Proposition 1.5 is equivalent to the following.
@@
Let $c>0$ be an absolute constant to be fixed shortly. If $C$ has radius at most $cN^{2}$, then $|S|\ll N^{O(1/\log\log N)}$ by Proposition 5.1. Thus, the proof is complete in this case.
@@
then, by [30, Observation 1], we have
@@
As the diameter of $[-N,N]^{2}$ is $\ll N$, the intersection of $C$ with $[-N,N]^{2}$ must be contained in an arc $A$ of $C$ of length $\ll N$. Therefore, the central angle $\alpha$ of $A$ in $C$ satisfies $\alpha r\ll N$, which gives
@@
$-Da^{2}=m^{2}-n^{2}=m^{2}-(n_{1}n_{2})^{2},$
it follows that both $a$ and $m$ are divisible by $n_{1}$, for otherwise $-D$ would be a quadratic residue modulo some prime dividing $n_{1}$, contradicting the definition of $n_{1}$. Setting $x=\frac{m}{n_{1}}$ and $y=\frac{a}{n_{1}}$, so that $x,y\in\mathbf{Z}$, we thus get
@@
The right-hand side above is well-known to be $\ll\tau(n_{2}^{2})$ (see, for example, [21, Chapter 11]). Since
@@ 6. Conclusion of the argument
By Chevalley’s theorem (see, for example, Exercise 3.19 of [18, Chapter I]), there exist a finite number of affine varieties $V_{1},V_{1}^{\prime},\dots,V_{n},V^{\prime}_{n}\subset\mathbf{C}^{2}$ such that
@@
Now we are ready to prove Theorem 1.1, from which Corollaries 1.3 and 1.6 immediately follow (the latter by invoking Propositions 1.4 and 1.5). In fact, we will prove an even stronger structure theorem that implies Theorem 1.1.
Theorem 6.3.
Let $S\subset[-N,N]^{2}$ be an integer distance set. Then either
$|S|\ll(\log{N})^{O(1)},$
or there exists a line or circle $C\subset\mathbf{R}^{2}$ such that
$|S\setminus C|\ll(\log\log{N})^{2}.$
Proof.
We follow the outline in Section 2. Let $S\subset[-N,N]^{2}$ be an integer distance set and let $k=k(N)\in\mathbf{N}$, with $k(N)\to\infty$ as $N\to\infty$ and $k(N)\ll\log\log{N}$, be a parameter to be chosen later. We may assume, without loss of generality, that $S$ contains the origin and at least $k$ other points, and, further, that there exists an $M\in\mathbf{N}$ and a squarefree $m\in\mathbf{N}$ such that
@@
of height $\ll N^{2}$. By Lemma 3.1, $\overline{X_{k}}$ is an irreducible surface of degree $2^{k}$. Thus, it follows from Theorem 2.2 and Lemma 6.1 that $S$ is contained in the union of $t\ll e^{O(k)}N^{3/2^{k/2+1}}$ irreducible affine curves $C_{1},\dots,C_{t}$, each of degree at most $s\ll e^{O(k)}N^{3/2^{k/2+1}}$.
For each $j=1,\dots,t$, either $C_{j}$ contains at least $(\deg{C_{j}})^{2}+1$ points in $S$, in which case $\left\{(x,y)\in\mathbf{C}^{2}:(x,y\sqrt{m})\in C_{j}\right\}$ is defined over $\mathbf{Q}$ by Lemma 6.2, or else $S\cap C_{j}$ contains at most $(\deg C_{j})^{2}\ll e^{O(k)}N^{3/2^{k/2}}$ points. It follows that all but $\ll e^{O(k)}N^{5/2^{k/2}}$ points of $S$ are contained in the union of $t^{\prime}\ll e^{O(k)}N^{3/2^{k/2+1}}$ irreducible affine curves $\tilde{C}_{1},\dots,\tilde{C}_{t^{\prime}}$ of degree at most $s$, where each
@@
Let $1\leq j\leq t^{\prime}$. Assume first that $\tilde{C}_{j}$ is not a line or circle. We will now bound $|S\cap\tilde{C}_{j}|$. We may assume, without loss of generality (by applying a translation), that $S\cap\tilde{C}_{j}$ contains the origin and that the origin is not a singular point of $\tilde{C}^{\prime}_{j}$. Recall that an irreducible affine curve in $\mathbf{C}^{2}$ of degree $d$ can have at most $\frac{(d-1)(d-2)}{2}$ singular points (see, for example, Exercise 3 of [28, Chapter 3 §2]). Denoting the set of singular points of $\tilde{C}_{j}^{\prime}$ by $\mathcal{S}$, either $S\cap\tilde{C}_{j}$ contains at most
@@
$\{(a_{n},b_{n}):n=1,\dots,k\}\cap\bigcap_{l=0}^{\deg{\tilde{C}_{j}^{\prime}}-2}V(\tilde{R}_{l,+},\tilde{R}_{l,-})=\emptyset,$ (6.4)
@@
each of which consists exactly of the points $(a,b)\in\tilde{C}_{j}^{\prime}$ with $T_{\pm}(a,b)\in\tilde{C}^{\prime}_{j}+c_{n,\pm}$, where $T_{\pm}$ is a linear transformation that maps $\mathbf{C}^{2}$ to the line ±$\ell_{\pm}=V(x\pm i\sqrt{m})$ and $c_{n,\pm}\in\mathbf{C}^{2}$. For each choice of $1\leq n\leq k^{\prime}$ and sign $+$ or --, there are at most $\deg\tilde{C}_{j}^{\prime}$ points $z$ in ±$\ell_{\pm}\cap(\tilde{C}_{j}^{\prime}+c_{n,\pm})$ by Bézout’s inequality (Lemma 3.6), and, for each such $z$, there are at most $\deg\tilde{C}_{j}^{\prime}$ points $(a,b)\in\tilde{C}_{j}^{\prime}$ with $T_{\pm}(a,b)=z$, again by Bézout’s inequality (since $T_{\pm}(a,b)=z$ gives the equation of a line in the variables $a$ and $b$, while $\tilde{C}_{j}^{\prime}$ is not a line). Therefore, each of the above intersections can contain at most $(\deg{\tilde{C}^{\prime}_{j}})^{2}$ points; by summing
@@
in $\mathbf{C}^{k+2}$. Every point in $S\cap\tilde{C}_{j}$ corresponds to a rational point on $\overline{\mathcal{C}_{k}}$ of height $\ll N^{2}$. By Lemma 4.2, $\overline{\mathcal{C}_{k}}$ is an irreducible curve of degree at least $2^{k}$ and at most $2^{k}s$, so that, by Theorem 2.3, $|S\cap\tilde{C}_{j}|\ll e^{O(k)}N^{O(2^{-k})}$ whenever $S\cap\tilde{C}_{j}$ contains at least (6.2) nonsingular points of $\tilde{C}_{j}$. If $S\cap\tilde{C}_{j}$ contains fewer than (6.2) nonsingular points of $\tilde{C}_{j}$, then $|S\cap\tilde{C}_{j}|\ll k^{2}s^{2}\ll e^{O(k)}N^{O(2^{-k})}$. Thus, either way, we have that
@@
Now assume that $\tilde{C}_{j}$ is a line or circle. To bound $|S\cap\tilde{C}_{j}|$, we proceed as in the previous case, with a few modifications. We may again assume, without loss of generality, that $S\cap\tilde{C_{j}}$ contains the origin. Either $S$ contains at most $2+2k(k-1)$ points outside of $\tilde{C}_{j}$, or else there exist $k$ distinct points
@@
we conclude that as long as there are at least $2+2k(k-1)$ points outside of $\tilde{C}_{j}$ in $S$, we can find a suitable choice of $\tilde{P}_{1}^{\prime}=(a_{1}^{\prime},\sqrt{m}b_{1}^{\prime}),\dots,\tilde{P}_{k}^{\prime}=(a_{k}^{\prime},\sqrt{m}b_{k}^{\prime})$. We then set $P^{\prime}_{j}=(a_{j}^{\prime},b_{j}^{\prime})$ for $j=1,\dots,k$ and, again, consider the affine variety
@@
By Lemma 4.3, $\overline{\mathcal{C}}_{k}$ is an irreducible curve of degree between $2^{k}$ and $2^{k+1}$. Therefore, by Theorem 2.3, $|S\cap\tilde{C}_{j}|\ll e^{O(k)}N^{O(2^{-k})}$ whenever $S\setminus\tilde{C}_{j}$ contains at least $2+2k(k-1)$ points.
Thus, if, for all $1\leq j\leq t^{\prime}$, we have that either $\tilde{C}_{j}$ is not a line or circle or $\tilde{C}_{j}$ is a line or circle and $S\setminus\tilde{C}_{j}$ contains at least $2+2k(k-1)$ points, then
$|S|\ll t^{\prime}e^{O(k)}N^{O(2^{-k})}\ll e^{O(k)}N^{O(2^{-k})}.$
@@
Finally, if, for some $1\leq j\leq t^{\prime}$, $\tilde{C}_{j}$ is a line or circle and $S\setminus\tilde{C}_{j}$ contains fewer than $2+2k(k-1)$ points, we certainly have that all but at most $\ll(\log\log{N})^{2}$ points of $S$ lie on a single line or circle. This completes the proof of the theorem. ∎

### v3

@@ 5.1. Lines
Proposition 1.5 was essentially proved in [31, Section 5]. However, since some of the terms used in [31] are nonstandard and [31, Theorem 4] is phrased in a manner that is not equivalent to Proposition 1.5, we will present a detailed proof of Proposition 1.5 here.
@@
$\left|\left\{Db_{i}+D(q_{1}-m_{i})\colon i=1,\dots,K\right\}\right|\asymp K.$
@@
using the standard estimate $\tau(n)\ll n^{O(1/\log\log n)}$ (see, for example, [20, Theorem 315]), which gives the claim. ∎
@@ 5.2. Circles
Recall that Proposition 1.6 controls the size of an integer distance set in $N,N]^{2}$ that lives on an arbitrary circle. It turns out that the bulk of the argument lies in tackling the case where the circle is fully contained in $N,N]^{2}$; the general case easily follows. In particular, for any positive real number $r$, define $c(r)$ to be the maximum size of an integer distance set contained in the circle of radius $r$ centered at the origin. Then, this special case of Proposition 1.6 is equivalent to the following.
@@
Let $c>0$ be an absolute constant to be fixed shortly. Since $N$ is assumed to be large, if $C$ has radius at most $cN^{2}$, then $|S|\ll N^{O(1/\log\log N)}$ by Proposition 5.1. Thus, the proof is complete in this case.
@@
then, by [38, Observation 1], we have
@@
As the diameter of $N,N]^{2}$ is $\ll N$, the intersection of $C$ with $N,N]^{2}$ must be contained in an arc $A$ of $C$ of length $\ll N$. Therefore, the central angle $\alpha$ of $A$ in $C$ satisfies $\alpha r\ll N$, which gives
@@
$Da^{2}=m^{2}-n^{2}=m^{2}-(n_{1}n_{2})^{2},$
it follows that both $a$ and $m$ are divisible by $n_{1}$, for otherwise D-D would be a quadratic residue modulo some prime dividing $n_{1}$, contradicting the definition of $n_{1}$. Setting $x=\frac{m}{n_{1}}$ and $y=\frac{a}{n_{1}}$, so that $x,y\in\mathbf{Z}$, we thus get
@@
The right-hand side above is well-known to be $\ll\tau(n_{2}^{2})$ (see, for example, [28, Chapter 11]). Since
@@ 6. Conclusion of the argument
By Chevalley’s theorem (see, for example, Exercise 3.19 of [21, Chapter I]), there exist a finite number of affine varieties $V_{1},V_{1}^{\prime},\dots,V_{n},V^{\prime}_{n}\subset\mathbf{C}^{2}$ such that
@@
Now we are ready to prove Theorem 1.1, from which Corollaries 1.3 and 1.7 immediately follow (the latter by invoking Propositions 1.5 and 1.6).
Proof od Theorem 1.1.
We follow the outline in Section 2. We will assume that $N$ is larger than any specified absolute constant, since the results are trivial for bounded $N$. Let $S\subset[-N,N]^{2}$ be an integer distance set and let $k=k(N)\in\mathbf{N}$, with $k(N)\to\infty$ as $N\to\infty$ and $k(N)\ll\log\log{N}$, be a parameter to be chosen later. We may assume, without loss of generality, that $S$ contains the origin and at least $k$ other points, and, further, that there exists an $M\in\mathbf{N}$ and a squarefree $m\in\mathbf{N}$ such that
@@
of height $\ll N^{2}$. By Lemma 3.1, $\overline{X_{k}}$ is an irreducible surface of degree $2^{k}$. Thus, it follows from Theorem 2.2 and Lemma 6.1 that $S$ is contained in the union of $\tau\ll e^{O(k)}N^{3/2^{k/2+1}}$ irreducible affine curves $C_{1},\dots,C_{\tau}$, each of degree at most $\sigma\ll e^{O(k)}N^{3/2^{k/2+1}}$.
For each $j=1,\dots,\tau$, either $C_{j}$ contains at least $(\deg{C_{j}})^{2}+1$ points in $S$, in which case $\left\{(x,y)\in\mathbf{C}^{2}:(x,y\sqrt{m})\in C_{j}\right\}$ is defined over $\mathbf{Q}$ by Lemma 6.2, or else $S\cap C_{j}$ contains at most $(\deg C_{j})^{2}\ll e^{O(k)}N^{3/2^{k/2}}$ points. It follows that all but $\ll e^{O(k)}N^{5/2^{k/2}}$ points of $S$ are contained in the union of $\tau^{\prime}\ll e^{O(k)}N^{3/2^{k/2+1}}$ irreducible affine curves $\tilde{C}_{1},\dots,\tilde{C}_{\tau^{\prime}}$ of degree at most $\sigma$, where each
@@
Let $1\leq j\leq\tau^{\prime}$. Assume first that $\tilde{C}_{j}$ is not a line or circle. We will now bound $|S\cap\tilde{C}_{j}|$. We may assume, without loss of generality (by applying a translation), that $S\cap\tilde{C}_{j}$ contains the origin and that the origin is not a singular point of $\tilde{C}^{\prime}_{j}$. Recall that an irreducible affine curve in $\mathbf{C}^{2}$ of degree $d$ can have at most $\frac{(d-1)(d-2)}{2}$ singular points (see, for example, Exercise 3 of [36, Chapter 3 §2]). Denoting the set of singular points of $\tilde{C}_{j}^{\prime}$ by $\mathcal{S}$, either $S\cap\tilde{C}_{j}$ contains at most
@@
$\left\{(a_{n},b_{n}):n=1,\dots,k\right\}\cap\bigcap_{l=0}^{\deg{\tilde{C}_{j}^{\prime}}-2}V(\tilde{R}_{l,+},\tilde{R}_{l,-})=\emptyset,$ (6.4)
@@
each of which consists exactly of the points $(a,b)\in\tilde{C}_{j}^{\prime}$ with $T_{\pm}(a,b)\in\tilde{C}^{\prime}_{j}+c_{n,\pm}$, where $T_{\pm}$ is a linear transformation that maps $\mathbf{C}^{2}$ to the line $\ell_{\pm}=V(x\pm i\sqrt{m})$ and $c_{n,\pm}\in\mathbf{C}^{2}$. For each choice of $1\leq n\leq k^{\prime}$ and sign ++ or -, there are at most $\deg\tilde{C}_{j}^{\prime}$ points $z$ in $\ell_{\pm}\cap(\tilde{C}_{j}^{\prime}+c_{n,\pm})$ by Bézout’s inequality (Lemma 3.6), and, for each such $z$, there are at most $\deg\tilde{C}_{j}^{\prime}$ points $(a,b)\in\tilde{C}_{j}^{\prime}$ with $T_{\pm}(a,b)=z$, again by Bézout’s inequality (since $T_{\pm}(a,b)=z$ gives the equation of a line in the variables $a$ and $b$, while $\tilde{C}_{j}^{\prime}$ is not a line). Therefore, each of the above intersections can contain at most $(\deg{\tilde{C}^{\prime}_{j}})^{2}$ points; by summing
@@
in $\mathbf{C}^{k+2}$. Every point in $S\cap\tilde{C}_{j}$ corresponds to a rational point on $\overline{\mathcal{C}_{k}}$ of height $\ll N^{2}$. By Lemma 4.2, $\overline{\mathcal{C}_{k}}$ is an irreducible curve of degree at least $2^{k}$ and at most $2^{k}\sigma$, so that, by Theorem 2.3, $|S\cap\tilde{C}_{j}|\ll e^{O(k)}N^{O(2^{-k})}$ whenever $S\cap\tilde{C}_{j}$ contains at least (6.2) nonsingular points of $\tilde{C}_{j}$. If $S\cap\tilde{C}_{j}$ contains fewer than (6.2) nonsingular points of $\tilde{C}_{j}$, then, as $N$ was chosen to be sufficiently large, $|S\cap\tilde{C}_{j}|\ll k^{2}\sigma^{2}\ll e^{O(k)}N^{O(2^{-k})}$. Thus, either way, we have that
@@
Now assume that $\tilde{C}_{j}$ is a line or circle. To bound $|S\cap\tilde{C}_{j}|$, we proceed as in the previous case, with a few modifications. We may again assume, without loss of generality, that $S\cap\tilde{C_{j}}$ contains the origin. Either $S$ contains at most 2+2k(k1)2+2k(k-1) points outside of $\tilde{C}_{j}$, or else there exist $k$ distinct points
@@
we conclude that as long as there are at least 2+2k(k1)2+2k(k-1) points outside of $\tilde{C}_{j}$ in $S$, we can find a suitable choice of $\tilde{P}_{1}^{\prime}=(a_{1}^{\prime},\sqrt{m}b_{1}^{\prime}),\dots,\tilde{P}_{k}^{\prime}=(a_{k}^{\prime},\sqrt{m}b_{k}^{\prime})$. We then set $P^{\prime}_{j}=(a_{j}^{\prime},b_{j}^{\prime})$ for $j=1,\dots,k$ and, again, consider the affine variety
@@
By Lemma 4.3, $\overline{\mathcal{C}}_{k}$ is an irreducible curve of degree between $2^{k}$ and $2^{k+1}$. Therefore, by Theorem 2.3, $|S\cap\tilde{C}_{j}|\ll e^{O(k)}N^{O(2^{-k})}$ whenever $S\setminus\tilde{C}_{j}$ contains at least 2+2k(k1)2+2k(k-1) points.
Thus, if, for all $1\leq j\leq\tau^{\prime}$, we have that either $\tilde{C}_{j}$ is not a line or circle or $\tilde{C}_{j}$ is a line or circle and $S\setminus\tilde{C}_{j}$ contains at least 2+2k(k1)2+2k(k-1) points, then
$|S|\ll\tau^{\prime}e^{O(k)}N^{O(2^{-k})}\ll e^{O(k)}N^{O(2^{-k})}.$
@@
Finally, if, for some $1\leq j\leq\tau^{\prime}$, $\tilde{C}_{j}$ is a line or circle and $S\setminus\tilde{C}_{j}$ contains fewer than 2+2k(k1)2+2k(k-1) points, as $N$ is chosen to be sufficiently large, we certainly have that all but at most $\ll(\log\log{N})^{2}$ points of $S$ lie on a single line or circle. This completes the proof of the theorem. ∎